### v2

# Equations of State, Thermodynamics, and Miscibility Curves for Jovian Planet and Giant Exoplanet Evolutionary Models

Abstract
The equation of state of hydrogen-helium (H-He) mixtures plays a vital role in the evolution and structure of gas giant planets and exoplanets. Recent equations of state that account for hydrogen-helium interactions, coupled with hydrogen-helium immiscibility curves, can now produce more physical evolutionary models, such as accounting for helium rain with greater fidelity than in the past. In this work, we present a set of tools for planetary evolution111All tables of thermodynamic quantities and derivatives are available at https://github.com/Rob685/hhe_eos_misc, along with a unified Python interface. Tutorials demonstrating the interface are also available in the repository. that provides a Python interface for existing tables of useful thermodynamic quantities, state-of-the-art H-He equations of state, and pressure-dependent H-He immiscibility curves. In particular, for a collection of independent variable choices, we provide scripts to calculate the variety of thermodynamic derivatives used to model convection and energy transport. These include the chemical potential derived from the internal energy, which is a modeling necessity in the presence of composition gradients when entropy is the other primary variable. Finally, an entropy-based convection formalism is presented and fully described that highlights the physical differences between adiabatic and isentropic interior models. This centralized resource is meant to facilitate both giant planet structural and evolutionary modeling and the entry of new research groups into the field of giant planet modeling.
1 Introduction
The structure and evolution of gas giant planets and brown dwarfs depend on the properties of hydrogen and helium at high pressures (Hubbard, 1970; Hubbard & Slattery, 1971; Hubbard & Dewitt, 1985; Burrows et al., 1997, 2001; Fortney & Hubbard, 2003, 2004; Militzer et al., 2008; Nettelmann et al., 2008, 2012, 2015; Hubbard & Militzer, 2016; Mankovich et al., 2016; Püstow et al., 2016; Miguel et al., 2016; Debras & Chabrier, 2019; Mankovich & Fortney, 2020; Debras et al., 2021a; Mankovich & Fuller, 2021; Howard et al., 2023), as well as their atmospheres (Burrows et al., 1997; Fortney et al., 2011; Burrows, 2014; Chen et al., 2023). The equation of state (EOS) most employed in the past was published by Saumon et al. (1995, SCvH95). This EOS incorporated the so-called “chemical picture” of the constituent species, obtained by minimizing the Helmholtz free energy of its constituents. This model estimated the gradual dissociation and ionization of hydrogen with increasing pressure, thus accounting for the increasing densities and even the predicted hydrogen first-order phase transition within Jupiter. Subsequent ab initio calculations by Militzer & Hubbard (2013, MH13), who included non-ideal interaction effects of hydrogen-helium mixtures and partial ionizations, were an important step forward. Models using updated EOSes by MH13 and Nettelmann et al. (2012) resulted in significant differences in the interior structures of Jupiter and Saturn (see Miguel et al., 2016; Militzer & Hubbard, 2013; Hubbard & Militzer, 2016; Miguel et al., 2022). Further complications arose when applying these EOSes to such modeling due to the phase separation of helium related to the metallization of hydrogen at high pressures; the presence of a higher-Z core; and the presence of a component of heavy elements in the gaseous mantle. The distribution of the latter is the subject of intense debate following the recent interpretations (Leconte & Chabrier, 2012, 2013; Vazan et al., 2016; Wahl et al., 2017; Folkner et al., 2017; Vazan et al., 2018; Stevenson, 2020; Helled et al., 2022; Militzer et al., 2022; Howard et al., 2023) of the Juno (Bolton et al., 2017a, b) gravity data. To this end, two primary areas of focus are an improved H-He EOS and a more accurate H-He phase separation curve.
The first area, the H-He EOS, demands ab initio calculations of relevant thermodynamic quantities. The computational costs have impeded the creation of a comprehensive tabular EOS for a wide range of hydrogen-helium mixtures, temperatures, and pressures. As a result, current EOSes designed to be employed in structural and evolutionary codes, such as those of SCvH95, Nettelmann et al. (2012); Becker et al. (2014), Chabrier et al. (2019, CMS19) and Mazevet et al. (2022, MLS22), rely on idealized mixing laws to compute the entropy of H-He mixtures. A new class of models includes non-ideal effects of the entropy of mixing (Militzer & Hubbard, 2013; Chabrier & Debras, 2021). These terms have recently been estimated by Howard & Guillot (2023, HG23) based on the calculations of MH13 and CD21, where MH13 calculated a H-He mixture EOS for a helium mass fraction () of 0.245. This resulted in estimates for the non-ideal corrections of the entropy and specific volume (density) for any helium mass fraction for the CMS19 and MLS22 EOSes. However, Howard & Guillot (2023) did not provide the corresponding correction for the internal energy.
In the second area, the H-He phase separation curves, Lorenzen et al. (2009, 2011), Morales et al. (2013), and Schöttler & Redmer (2018) have employed ab initio techniques to determine the phase separation curves of hydrogen and helium mixtures. On the experimental front, Brygoo et al. (2021) conducted cutting-edge experiments on H-He phase separation and found significant disagreements with the ab initio calculations. Moreover, those ab initio calculations are not accompanied by corresponding EOSes, and the CMS19 and MLS22 EOSes are not accompanied by miscibility curves. Therefore, the predominant hybrid approach used in planetary evolution calculations, combining EOSes and phase separation curves from different studies, is inconsistent. However, the utility of this approach cannot be denied, and the ease of use of both EOSes and miscibility curves is of great importance.
Between the different EOS and phase separation calculations by various authors, planetary modelers have a wide range of options. In this paper, we collect many major EOS and miscibility databases required for giant planet evolutionary and structural calculations and describe their strengths and weaknesses222We omit the H-He EOS of Nettelmann et al. (2012) and Becker et al. (2014) due to the lack of entropy calculations available in their tables. An integration technique to obtain the entropy for these tables depends on choices of initial thermodynamic values, making the comparisons between EOSes uncertain and subsequent derivative calculations subject to this uncertainty.. We invert each EOS on a dense grid and calculate all the necessary quantities and derivatives that play a role in the physics, structure, thermal evolution, and convection criteria within Jupiter and Saturn in particular, and giant exoplanets in general. Each EOS is calculated along pressure-temperature, density-temperature, and entropy-density axes that span the thermodynamic range of the interiors of gas giants. As discussed in Section 6, a proper treatment of convection requires the calculation of various derivatives. In addition to the various EOS tables, we furnish useful tables of the hydrogen-helium immiscibility curves of Lorenzen et al. (2009, 2011) and Schöttler & Redmer (2018). While these data are published and employed in some of the most recent planetary evolution calculations for Jupiter and Saturn (Nettelmann et al., 2015; Püstow et al., 2016; Mankovich et al., 2016; Mankovich & Fortney, 2020), smooth fits to these data are not easily come by. Though demixing temperatures for a broad range of helium abundances are what authors such as Lorenzen et al. (2009, 2011) and Schöttler & Redmer (2018) have provided, easily used fits to these data are not easily generated. We provide such smooth fits here. In this way, our work provides a more centralized and standardized approach by collecting the most recent EOS and H-He immiscibility data in a self-consistent format.
This paper is organized as follows. Section 2 discusses the various hydrogen-helium (H-He) EOSes, compares their properties, and discusses their applications. In Section 3, we provide an overview of the available miscibility calculations in the literature, compare the different demixing temperature calculations of each of the sources, compare the H-He demixing temperatures with different available H-He equations of state, and discuss what various authors have concluded concerning helium rain in Jupiter and Saturn based on these different demixing temperatures. Section 4 describes the different EOS table inversions, inversion techniques, interpolation methods, and thermodynamic table ranges. This section also describes the H-He EOS and miscibility code structure. In Section 5, we briefly discuss the thermodynamic consistency of the EOSes published by Saumon et al. (1995), citeChabrier2021, and Chabrier et al. (2019) (with the non-ideal corrections by Howard & Guillot, 2023)). In Section 6, we discuss the various derivatives necessary for convection and energy calculations in a planetary model. Section 7 showcases an example evolutionary calculation using the code APPLE (Sur et al., 2024) and the tables presented here. Finally, in Section 8, we provide a summary discussion and concluding remarks. In the appendices, we derive various of the relationships provided in the text.
2 Hydrogen and Helium Equations of State
Historically, the EOS of hydrogen and helium at high pressures has been divided into the “physical” picture and the “chemical” picture. In the physical picture, electrons and nuclei interact through Coulomb potentials, and the EOS is obtained by solving the resulting quantum many-body problem. In the chemical picture, atoms remain bound and interact through pair potentials. However, at ionization pressures and temperatures, electrons become unbound and delocalized. In the high temperature regime, electrons become unbound and the physical picture is tractable (Rogers, 1981; Rogers & Young, 1984). However, in the low-temperature and high-pressure regime relevant to gas giant planets, the physical picture could not easily handle partial ionization. The chemical picture presents a simplified model a physical model. In the past, the chemical picture of SCvH95 was the most commonly used H-He EOS. However, with improvements in computation rigorous physical (or “ab initio”) models have now been published. Most notably, MH13, CMS19, Chabrier & Debras (2021, CD21), and MLS22 have published EOSes obtained from ab initio calculations. All of these models rely on the additive volume law (AVL), since calculations for arbitrary hydrogen-helium mixtures remain intractable. This approximation was improved by HG23, who have calculated nonideal interaction corrections based on the MH13 results for the CMS19 and MLS22 EOSes. In this section, we discuss the properties of the historical SCvH95, the ab initio MH13, and the most c urrent CMS19 EOSes and compare their properties333The MLS22 EOS result is a hydrogen EOS table benchmarked with structural calculations of the interior of Jupiter inferred from the Juno data. Though we present this EOS and provide the same quantities and derivatives for it, we focus here on the CMS19 and SCvH95 EOSes..
2.1 Saumon, Chabrier, and van Horn (1995)
Saumon et al. (1995, SCvH95) employed the chemical picture and generated an EOS that covered a temperature range of and a density range of . They calculated pure hydrogen and helium tables and relied on the AVL, to handle a general helium mass fraction dependence. Intensive variables, such as temperature and pressure, must be uniform throughout a system at equilibrium. Extensive variables, such as volume and entropy, are additive for a system in equilibrium. Therefore, the AVL considers a combined system of equations of state and assumes they are in thermal and pressure equilibrium. AVL cannot account for interactions between hydrogen and helium, is not reliable in regions of partial ionization, and it does not account for phase separation (see Section 6 of SCvH95 for more detailed descriptions of its limitations). Despite these limitations, until recently, SCvH95 was the most used EOS in the study of low-mass stars, brown dwarfs, and giant planets (e.g., Burrows et al., 1995; Guillot et al., 1996; Saumon et al., 1996; Burrows et al., 1997; Chabrier & Baraffe, 1997; Baraffe et al., 1998; Burrows et al., 2001; Fortney & Hubbard, 2003, 2004; Fortney et al., 2011)
The additive volume law states that
where , , and are the mass densities of hydrogen, helium and mixtures, respectively, at constant pressure and temperature. This corresponds to
where , , and are the hydrogen, helium, and the ideal entropy of mixing, respectively. The SCvH95 EOS estimated entropy of mixing for each component of systems,
This generalizes to all ionization states in their EOS and was used to compute the ideal entropy of mixing for H-He mixtures (see equations 53–56 in SCvH95).
The SCvH95 EOS conveniently provides many of these derivatives, making it the standard for H-He EOSes in the warm-dense region of planetary interiors and low mass stars (Burrows et al., 1997, 2001). For all its strengths, the most challenging region for the SCvH95 EOS is that of partial ionization in the region of Mbar, which is relevant for gas giant interiors. The limits of the chemical paradigm, such as the ionization states of different chemical species, are overcome by ab initio simulations under the physical picture.
2.2 Militzer & Hubbard (2013)
The Militzer & Hubbard (2013, MH13) EOS spans a temperature range of K and a density range of g cm-3. As their model is a physical model, they avoid the limitations of the low-temperature and high-pressure regime inherent in the chemical model of SCvH95. They use Density Functional Theory Molecular Dynamics (DFT-MD), coupled with the thermodynamic integration (TDI) technique, to obtain the entropy directly through the difference in Helmholtz free energy between hydrogen and helium mixtures while using the Perdew-Burke-Ernzerhof (Perdew et al., 1996, PBE) exchange-correlation functional. To account for interactions between species and electrons via TDI, they introduced classical intramolecular pair potentials to match their DFT calculations as closely as possible (Militzer & Hubbard, 2013).
Since MH13 calculated their ab initio EOS only for , they used the AVL for arbitrary mixtures. Since the AVL requires another helium fraction, the EOS table can be combined with an ab initio helium EOS calculated by Militzer (2009). In principle any helium EOS can be used (e.g., see Miguel et al., 2016, who used the SCvH95 helium EOS). The AVL can be adjusted to interpolate between any given H-He mixture, and , by
where , and are the densities at and respectively. For a pure helium EOS and a mixture EOS, Equation 4 reduces to Equation 3 of Hubbard & Militzer (2016).
The MH13 EOS addressed many of the shortcomings of the SCvH95 EOS, but its lack of coverage in the shallow interior ( Mbar) prompted the use of other EOS prescriptions in such regions. To overcome these obstacles, recent work (see Mankovich & Fuller, 2021, for example) uses a combination of EOS tables, including that of MH13, for better coverage in the density, temperature, and pressure space and greater flexibility with helium fraction dependencies. Next, we discuss the most recent work by CMS19 and CD21 to obtain a comprehensive H-He EOS.
2.3 Chabrier, Mazevet, and Soubiran (2019)
Chabrier et al. (2019, CMS19) constructed pure hydrogen and helium EOS tables spanning wider density and temperature ranges, namely g cm-3 and K. For densities lower than 0.05 g cm-3, they use the SCvH95 hydrogen EOS, the hydrogen EOS of Caillabet et al. (2011) for g cm-3, and the EOS of Chabrier & Potekhin (1998) for densities greater than 10 g cm-3.
As does SCvH95, their EOS relies on the AVL for helium mixtures and includes only the ideal entropy of mixing. CD21 incorporate the non-ideal interaction effects of MH13 into the CMS19 EOS in a manner similar to Miguel et al. (2016). Whereas Miguel et al. (2016) did not account for the non-ideal entropy contributions from MH13, CD21 did. Using the AVL, CD21 obtained an “effective” hydrogen EOS from MH13 using the CMS19 helium EOS to obtain a mixture EOS that agrees with MH13 at (see Figure 9 of CD21).
To further address this shortcoming, Howard & Guillot (2023, HG23) extended the interaction correction to arbitrary by prescribing a non-ideal correction term of form
and requiring that these corrections agree with MH13 at and with CD21 at . The corrections, and , are assumed to be proportional to the products of hydrogen and helium mass fractions.
The CMS19 EOS with the HG23 corrections (CMS19+HG23) and the CD21 EOS are the most current development of an H-He EOS that spans all relevant thermodynamic space, helium fractions, and nonideal interactions. In Figure 1, we compare the CMS19+HG13, CD21, and MLS22+HG23 EOSes with the SCvH95, MH13, and CMS19 tables, illustrating the differences among them. The top panel mirrors 2 of Sur et al. (2024), originally inspired by HG23, and compares the EOSes at a helium fraction of and a specific entropy of baryon. The bottom panel shows the temperature differences due to the non-ideal interactions at different entropy values, as noted above, for each profile. The density deviations from SCvH95 increase with decreasing entropy, particularly between 0.5 and 7 Mbar. This pressure region is expected to exhibit H-He phase separation at a sufficiently low temperature (Stevenson, 1975; Stevenson & Salpeter, 1977a). Figure 1 also compares the EOS adiabats to a 1 MJ profile at 0.5 and 5 Gyrs using APPLE (Sur et al., 2024) with the CD21 EOS. A full example of a 1 MJ evolution with APPLE is discussed in Section 7. In the next section, we discuss the current understanding of H-He immiscibility in this context.
3 Hydrogen and Helium Immiscibility
Early evolutionary calculations for Saturn (Pollack & Graboske, 1976; Pollack et al., 1977) could not reproduce its current heat flux. To account for this apparent excess heat, Stevenson (1975) proposed the formation of He-rich droplets that rain into the interior (also see Stevenson & Salpeter, 1977a). One of the first attempts to capture and apply the demixing of hydrogen and helium as a function of helium abundance was done by Hubbard & Dewitt (1985), who applied Monte Carlo simulations for a fully ionized hydrogen and helium plasma. The low-pressure regime was studied by Schouten et al. (1991) via Monte Carlo methods using pair potentials for interactions. Pfaffenzeller et al. (1995) calculated the demixing temperatures using molecular dynamics (MD) and found a demixing temperature-pressure dependence opposite that of Hubbard & Dewitt (1985). However, the demixing temperatures remained too cold for Saturn to ever reach the immiscibility regions (Pfaffenzeller et al., 1995; Fortney & Hubbard, 2003). Fortney & Hubbard (2003) created an ad hoc phase diagram, adjusting the demixing temperatures of Pfaffenzeller et al. (1995), to match the present effective temperature of Saturn. More recently, H-He phase diagrams have been calculated by Lorenzen et al. (2009, 2011, L0911), Morales et al. (2009, 2013), and Schöttler & Redmer (2018, SR18). These ab initio phase diagrams are discussed in the next sections.
3.1 Lorenzen et al. (2009, 2011)
Lorenzen et al. (2009, L0911) conducted the first ab initio immiscibility gap model. They used DFT-MD with a PBE exchange-correlation functional and pair potentials for hydrogen and helium. They calculated the demixing temperatures for pressures between 4 Mbar and 24 Mbar and expanded their model in Lorenzen et al. (2011) to include 1 and 2 Mbar. In all cases, they relied on linear mixing to calculate the enthalpy of mixing to then infer the change in the Gibbs free energy, . This was then minimized to constrain the H-He miscibility gaps. They linearized the change in the enthalpy with
where is the number fraction of helium. They used the relation,
where ] is the change in the ideal entropy of mixing.
With a helium abundance-dependent phase diagram, one can interpolate between the isobars and invert it to yield the demixing temperature profiles. The right panel of Figure 2 shows the helium dependence of the demixing temperatures for each isobar calculated by L0911. Temperatures below each of these lines indicate regions where the hydrogen and helium phase separate. In Figure 3, the SCvH95 (solid lines) and CMS19+HG23 (dotted lines) temperature profiles are calculated for a constant helium fraction of and a specific entropy of baryon, believed to be the present-day entropy of an adiabatic Jupiter as inferred by the Galileo probe entry (Seiff et al., 1997, 1998). The bottom panel shows the demixing profiles of L0911. For example, since both EOS profiles intercept the demixing curve (in red), it is expected that demixing occurs between 1 Mbar and 2 Mbar for SCvH and 1 Mbar and 5 Mbar for the CMS19+HG3 temperature profile.
It has been suggested by both Püstow et al. (2016) and Mankovich et al. (2016) that the unperturbed phase diagrams of L0911 are unlikely due to their high demixing temperatures. Püstow et al. (2016) found that to obtain for Saturn an atmospheric helium mass fraction between and a shift of the L0911 miscibility curve needed to range from K to K, respectively. For Jupiter, Mankovich et al. (2016) found that to fit the observed value of (Von Zahn et al., 1998) this median shift of K. In the next section, we discuss the work of Morales et al. (2009, 2013) and their inclusion of non-ideal interactions in the phase diagram.
3.2 Morales et al. (2009, 2013)
Morales et al. (2009) constructed the first H-He phase diagram that included non-ideal interaction effects in the entropy of mixing, following a similar methodology to that of Militzer (2009) and MH13. Morales et al. (2009) used the Born-Oppenheimer MD coupled with the PBE exchange-correlation functional, and used TDI to infer the entropy using the Helmholtz free energy difference (also used by MH13). Morales et al. (2013) expanded their work to pressures below 0.4 Mbar, and claimed that the linear mixing assumption leads to significant disagreements with their direct calculation methods at high helium mass fractions and high pressures.
This work led to colder demixing temperatures compared to L0911. While Jupiter had a small miscibility gap between Mbar and Mbar with the L0911 phase diagram, the Morales demixing temperature results were too cold to allow Jupiter any miscibility gap (Morales et al., 2013). For consistency, Hubbard & Militzer (2016) used the Morales phase diagram, since both the EOS and the demixing temperatures were calculated in a similar fashion. Hubbard & Militzer (2016) then claimed that helium rain occurs within Jupiter at an entropy value of baryon. However, as Mankovich & Fortney (2020) also discuss, the published demixing temperatures are restricted to only the solar helium number fraction of . Given that helium is expected to redistribute inside any gas giant planet, a faithful model of the interiors of these planets that takes into account helium rain must be a function of helium abundance. In the following section, we discuss the work of Schöttler & Redmer (2018), which used ab initio methods, including non-ideal effects, but also as a function of helium abundance.
3.3 Schöttler & Redmer (2018)
Schöttler & Redmer (2018, SR18) performed ab initio simulations using a van der Waals exchange-correlation functional instead of PBE, claiming that their method better captured the Gibbs free energy compared to previously used methods. By calculating the non-ideal effects in the entropy, SR18 calculated the demixing temperatures shown on the left panel of Figure 2. It is clear that each isobar from SR18 is significantly colder than those of L0911. The 1 Mbar demixing region is significantly smaller for the van der Waals DFT result. Qualitatively, whereas L0911 showed enclosed regions of miscibility at low pressures, SR18 show flat and open regions at those pressures, leading to even colder demixing temperatures. The top panel of Figure 3 shows the demixing temperature profiles with colors corresponding to those of L0911 in the bottom panel. Note that for very low helium number fractions, the temperatures are comparable, but remain significantly lower for higher helium abundances.
When the demixing profiles are compared in the abundance of 444This is the primordial solar helium abundance value obtained from from Bahcall et al. (2006), where , ., the abundance for which Morales calculated their miscibility curves, the L0911 and Morales curves are hotter than the SR18 profiles by K and K, respectively. The top panel of Figure 4 shows these demixing temperature profiles (colored lines) compared to adiabats for the SCvH95 and CMS19+HG23 equations of state. If the demixing temperatures of SR18 are used to determine miscibility gaps in gas giant planets, then these will lead to narrower H-He immiscibility regions compared to L0911. The bottom panel of Figure 4 shows decreasing entropy adiabats for both EOSes. Once the adiabat intercepts the miscibility curve, the regions highlighted in red undergo H-He phase separation. For a constant helium fraction of , this occurs at baryon for the CMS19+HG23 EOS and roughly at baryon for the SCvH EOS. We show the L0911 and SR18 miscibility gaps as a function of entropy and helium mass fraction in Figure 5. The helium rain regions grow as entropy and helium mass fraction decrease. As helium mass fraction increases, immiscibility regions open only at lower entropies. Figure 5 illustrates the difference of each miscibility diagram, where L0911 can achieve miscibility gaps at higher entropies than SR18. Since the SR18 curves are significantly cooler, they require lower entropies. Because of this, Mankovich & Fortney (2020) found that if helium rain is responsible for the excess heat of Saturn, the H-He demixing temperatures of the SR18 miscibility curves need to be shifted by K. Recently, Brygoo et al. (2021) empirically measured a miscibility curve for , highlighted in red in the top panel of Figure 4. While they found a positive upward shift in the temperature from the SR18 curve, the shift is K, which is far larger than the shift predicted by Mankovich & Fortney (2020).
The SR18 H-He miscibility data have several advantages over the L0911 and Morales data. Its advantage over the L0911 phase diagram is its inclusion of non-ideal entropy of mixing, while both the SR18 and L0911 data have an advantage over the Morales phase diagram in its provision of demixing temperatures for an arbitrary helium mass fraction. We make the L0911 and SR18, and the Brygoo et al. (2021) experimental immiscibility curves conveniently accessible to use for giant planet evolutionary and structural calculations555All smooth miscibility curves are provided here: https://github.com/Rob685/hhe_eos_misc/tree/main/misc. These H-He miscibility prescriptions are allowed to follow the pressure, temperature, and helium abundance of a given planetary interior profile. These adapt to the conditions of the interior rather than remaining static; i.e., as the pressure and temperature of the planet interior changes, so does the H-He immiscibility curve. Note that these miscibility curves were originally calculated using pure H-He mixtures. Therefore, these curves do not account for alterations of the planetary profiles due to metals. Moreover, the H-He EOSes used to calculate the immiscibility curves by SR18 and L0911 differ from those provided here.
4 EOS & Miscibility tables and methods
This work provides three additional inverted tables in addition to the original tables published by SCvH95, MH13, CMS19, CD21, and MLS22. As shown in Table 1, the tables computed here have the following independent thermodynamic variable pairs: ; ; and . These tables have been computed across values of and . All table ranges are described below. While SCvH95 and CMS19 have published tables in both pressure-temperature and density-temperature, EOS mixtures of and using the AVL (Eqs. 5 and 6) can be performed only when combining entropy, density, and internal energy (e.g., Saumon et al., 1995). Tables with as independent variables require this density be the density of the mixture and the temperature of the mixture. Likewise, tables with independent variables require that the particular independent entropy and density be the mixture quantities, and not those of the original H and He tables.
To meet this requirement, numerical inversions are performed to obtain the appropriate independent thermodynamic variables; i.e., if an EOS call requires the temperature as a function of entropy, pressure, Y, and Z, then the entropy and pressure should be that of the H-He-metal mixture, and the appropriate temperature should be returned. We use a modified Powell method from MINPACK (More et al., 1980) as applied in SciPy (Virtanen et al., 2020)666The multi-variable inversions were performed using Scipy’s root application of MINPACK’s hybr FORTRAN routines. using an ideal H-He EOS as an initial guess. Such root finding methods allows us to recover original table values, and values between the original table points, to fractional errors between to . All original and subsequent inverted tables are interpolated using piecewise multi-linear methods (Weiser & Zarantonello, 1988; Virtanen et al., 2020).777The 4-dimensional linear interpolations are performed via Scipy.
4.1 EOS table ranges
We list all the thermodynamic variables for each table in Table 1 and the units are specified in Table 2. Each H-He EOS is combined with the pure water AQUA table (Haldemann et al., 2020). For brevity, each H-He-Z mixture table name is abbreviated with +AQUA to indicate a H-He-Z mixture. The () CMS19+HG23+AQUA and CD21+AQUA tables range from kbaryon to kbaryon, and the SCvH95+AQUA and MLS22+HG23+AQUA () tables range from kbaryon to kbaryon. All tables range from to and from to . The () tables cover the original temperature ranges, except for temperatures above K. 888All the EOS tables can be extrapolated linearly from the bounds indicated, but extrapolation is not advised. The H-He EOS tables are square in shape and are designed to cover the thermodynamic space of gas-giant planet evolution, but are not yet designed to cover temperatures above Kelvin and densities above 100 g cm-3.
4.2 H-He and Miscibility Package description
Each miscibility curve has its own module:
L0911: misc_lor.py
SR18: misc_sch.py
Brygoo et al. (2021): misc_brygoo.py .
Each of these modules contains T in units of K and P in Mbar, where Tm is the demixing temperature of each curve as a function of pressure and helium mass fraction and Pm is its inversion. Each of these functions has an option to offset the demixing temperatures by an amount in Kelvin. These functions were used to compute and show the miscibility curves in Figures 3 and 4.
A wrapper module called misc.py calls each of these files and offers users a method to calculate intercepting points between a choice of a miscibility curve and any temperature-pressure profile to generate a miscibility region. This function is used in APPLE’s miscibility scheme 1 (Sur et al., 2024). The misc.py module also contains a way to calculate Y, or the minimum helium abundance required for H-He immiscibility for a given pressure and temperature, as well as its derivative with respect to temperature. The Ym function is used in APPLE’s miscibility scheme 2.
As with the miscibility curves, and as described in the EOS repository tutorials999https://github.com/Rob685/hhe_eos_misc/blob/main/eos_tutorial.ipynb, each H-He EOS shown in Figure 1 has its own EOS module:
SCvH95: scvh_eos.py
MH13: mh13_eos.py
CMS19 (with or without HG23): cms_eos.py
CD21: cd_eos.py
MLS22: mls_eos.py
as well as the Z tables:
Water (AQUA): aqua_eos.py
Post-perovskite: ppv_eos.py
Iron: fe_eos.py
Serpentine (ANEOS): serpentine_eos.py .
For H-He-Z mixtures, we include a module called mixtures_eos.py that combines the user’s choice of H-He EOS with the user’s choice of Z EOS. The post-perovskite and iron EOS tables are provided by Zhang & Rogers (2022) via private communication, and the serpentine EOS from the analytical equation of state (ANEOS Thompson, 1990). All of these files contain the pre-calculated inverted tables described in this section and in Table 1, the inversion functions used for each table, and the derivatives.
4.3 Derivatives
All EOS modules described and listed above offer their own derivatives, but, as described in our repository derivatives tutorial101010https://github.com/Rob685/hhe_eos_misc/blob/main/eos_derivatives_tutorial.ipynb, mixtures_eos.py houses all derivatives in Table 1 for each H-He-Z mixture. For each derivative, the respective tables shown in the column headers of Table 1 are called to perform either a first order or second order finite difference. The user reserves control over the finite-difference deltas. Examples of a few derivatives used in energy transport are shown in Figure 7 and Figure 8 using fractional differences of along constant entropies and pressures for the purpose of differentiation. The derivatives shown in Figure 8 are used in APPLE to compute the Schwarzchild criteria () and the Ledoux criteria ). We discuss our convection convention in Section 6. Each of the files listed in this section can be accessed by following the installation instructions in the repository page.
5 Thermodynamic Consistency
The thermodynamic consistency for any EOS requires that its temperature, pressure, and internal energy be in agreement with the first law of thermodynamics, namely
and,
In Figure 6, we show regions of thermodynamic space relevant to planetary evolution with shaded regions where each EOS deviates from consistency. The top row shows the thermodynamic consistency of the CMS19+HG23 EOS, the middle row that of the CD21 EOS, while the bottom row shows the same for the SCvH95 EOS. Thermodynamic consistency is defined here as the difference between Equations 9 and 10 and the EOS table values over a 2-D grid of entropy and density values. In the CMS19+HG23 EOS, the metallic hydrogen phase transition is less thermodynamically consistent (denoted by darker colors), with overall errors of in the temperature. In the SCvH95 EOS, less thermodynamically consistent regions are found at higher entropies, due to interpolation techniques used by SCvH95 across the molecular/metallic transition and the plasma phase transition, with overall errors of in the temperature. These inconsistencies are in approximate agreement with the thermodynamic (in)consistency seen by Becker et al. (2014), and they imply that energy conservation of planetary evolution models can only be as good as their EOSes allow. Crucially, the internal energy of the CMS19+HG23 EOS does not yet account for non-ideal interactions. This inconsistency should be partially resolved once these internal energy terms become available. On the other hand, the internal energy values of the CD21 EOS already account for these non-ideal interactions, albeit at a constant helium fraction. Unlike the SCvH95 EOS, the CD21 does not interpolate across the plasma phase transition region. This accounts for the thermodynamic consistency observed in the CD21 EOS of under . Thermodynamic consistency must be considered when accounting for energy transport and overall energy conservation in evolutionary calculations.
6 Convection Criteria and The Energy Equation
In the standard mixing-length convection theory (Weiss et al., 2004; Kippenhahn & Weigert, 1990; Kippenhahn et al., 2012; Joyce & Tayar, 2023), the associated gedanken thought experiment, a fluid element is displaced and kept isolated at constant pressure with its new surroundings. Isolation in this context means the fluid element maintains a constant entropy and composition. Various structural and thermodynamic derivatives emerge to be important in the context of such theories of convection and convective transport. For simplicity, we evaluate these criteria assuming that the only compositional variation is in the helium mass fraction, , but these expressions can easily be generalized to arbitrarily complex compositional variation. The derivatives in question include:
and
An exhaustive table of the available values and derivatives for each H-He EOSes is shown in Table 1. Units used in our code are listed in Table 2. In Figure 7, we present , , , and . The entropy values selected for and encompass a range relevant to gas giant planet evolution, and the pressures selected for , and encompass the interior regions relevant for H-He phase separation.
Note. — The thermodynamic quantities and derivatives are also available for pure water (AQUA), post-perovskite, serpentine, and iron, along with the provided “metal” mixture tables111111All tables are accessible at our EOS repository: https://github.com/Rob685/hhe_eos_misc. The logarithmic quantites are log base 10, and the units are tabulated in Table 2. The original CMS19 tables were provided by their authors in the bases and . To ensure consistency, we used the original pressure-temperature combined with the HG23 components to obtain new tables. All quantities and derivatives of the CMS19 and MLS22 tables contain these non-ideal corrections.
Quantity | Units |
(independent) | kbaryon |
(dependent) | erg g-1 K-1 |
g cm-3 | |
dyn cm-2 | |
K | |
erg g-1 |
Note. — Standard units used throughout our EOS module. The EOS module provides thermodynamic quantites as a function of entropy; the independent entropy is expressed in units of kbaryon. If entropy is the dependent variable, our code will output the entropy in cgs units. We do not provide the internal energy as an independent variable.
6.1 Convection
The relative density of this isolated fluid compared to that of its surroundings is the criterion for convective stability. In the absence of a composition gradient, the Schwarzschild criterion for convective stability is
In the presence of a composition gradient, the Ledoux criterion (Ledoux, 1947) for stability must be used instead,
where is often defined as (Unno et al., 1989)
The composition gradient term, , can also be expressed in terms of entropy and composition gradients, derived in Appendix A, as
where
such that and is the pressure scale height, . The split into and is particularly useful. To whit, the superadiabaticity, , can be shown to be
so the Schwarzschild criterion for convective stability can then be expressed as
Recall that in the absence of a composition gradient, marginal Schwarzschild stability (i.e., the transition from stability to instability) corresponds to isentropy (). However, Eq. 28 shows that in the presence of a composition gradient, the transition from stability to instability occurs for a non-zero entropy gradient (even if ).
Similarly,
so that the Ledoux criterion for convective stability can be expressed as
In cases where a fluid is stable against convection according to the Ledoux criterion, but Schwarzschild unstable, semiconvection can arise in the form of double diffusive convection (see Walin, 1964; Kato, 1966; Stevenson, 1985; Rosenblum et al., 2011; Leconte & Chabrier, 2012; Mirouh et al., 2012; Wood et al., 2013; Leconte & Chabrier, 2013; Nettelmann et al., 2015; Moll et al., 2016; Garaud, 2018, and references therein). In lieu of any detailed understanding of energy transport in semiconvective regions (Debras et al., 2021b; Mankovich & Fortney, 2020; Mankovich et al., 2016; Nettelmann et al., 2015; Leconte & Chabrier, 2012), recent planetary models have included a density ratio , which describes the degree to which temperature and composition stratifications contribute to the density stratification to trigger convection. This parameter is defined as (see Rosenblum et al., 2011; Mirouh et al., 2012, for discussions)
Eq. 22 can be written with as
The parameter can be taken as a free parameter that stretches from 0 (Schwarzschild) to 1 (Ledoux). In terms of entropy, the criterion becomes
If , then Eq. 33 reduces to the Schwarzschild criterion, Eq. 28. If , then Eq. 33 reduces to the Ledoux criterion, Eq. 30, modulo the factor.
In the repository, we provide these derivatives in functional form. Figure 8 shows the sign changes of the relevant derivatives for the convection criteria at typical values in the interior of gas giants. Note that when the composition gradient () is negative, a negative ensures that when the Schwarzschild convection stability condition is violated, the resulting entropy gradient is positive. This also holds for the Ledoux condition against convection, but for instead of . A sign change (see Figure 8) of at interior conditions indicates that the reverse occurs; namely, the entropy profile gradient becomes negative when the Ledoux convection stability condition is violated in these regions. The derivative at constant () increases with decreasing entropy for densities g cm-3. Note also that temperatures here and for realistic gas giants are significantly higher than the hydrogen solidus line found in the CMS19 EOS. The SCvH95 EOS (dashed lines) also exhibits this behavior. This is in contrast with the Schwarzchild derivative, , which has no sign change.
We also provide the expressions to compute the Brunt-Väisälä frequency (), which is the bouyant oscillation frequency in a stable stratified fluid. This frequency is defined as (Lattimer & Mazurek, 1981; Unno et al., 1989; Saio, 1993; Passamonti et al., 2009; Fuller et al., 2016; Mankovich & Fuller, 2021, also shown in Appendix A.):
which can be shown121212in Appendix A to be equal to
6.2 Latent Heat Correction Due to Helium Phase Separation
We note that the energy source that is the gravitational settling of helium should exceed by a large margin the energy source that is the latent heat of helium phase separation and that all current evolutionary codes properly handle the former. Though we emphasize that the Ledoux condition (Eqs 22 and 30) assumes that the fluid mixtures are fully miscible, something not expected in a multi-phase context, one can approximately generalize the Ledoux condition to account for latent heat evolution (though in a miscible context) by employing what is done in the meteorological literature (Stevens, 2005) vis à vis wet versus dry adiabats and by employing the generalization found in Tremblin et al. (2019). The result is a multiplication factor on (Eqs. 21 and 22) of the form:
where is the latent heat per mass and Y is the helium abundance at which immiscibility ensues at a given and . Assuming the latent heat estimate of Stevenson & Salpeter (1977b) of order to per helium nucleus, the magnitude of this correction at the immiscibility curve is generally close to one, ranging from 0.93 to 0.99 for a broad range of temperature/pressure points and miscibility models. For the helium rain regions shown in Figure 9, it is indeed close to one. Above the immiscibility curve this correction is one and far below it one must have a better sense of the variation in the latent heat as the helium abundance associated with the phase partitioning varies. However, until complete numerical data on the latent heats under all thermodynamic and compositional circumstances are available, the helium phase-separation latent heat effect on a generalized Ledoux criterion must, for all intents and purposes, be considered unknown.
6.3 The energy equation
The thermodynamic energy equation can be given by the first law of thermodynamics. If rendered in the independent thermodynamic basis variables (), this yields
for chemical components. It is the left-hand-side of this equation that appears in the energy equation in stellar/planetary evolution and that is related to the energy sources, sinks, and the divergence of the energy flux. If the thermodynamic basis variables are (), then a knowledge of allows one to eschew explicit mention of .
However, if we employ the entropy formulation and include only the dependence upon (the helium mass fraction), the specific energy (energy per mass) can be written as
We note that is not the helium chemical potential but the weighted sum emerging from the right-hand side of Eq. 38. This can be easily expanded to include metals () specifically by adding a term, but is in fact already implicit in Eq. 38. The composition then is included in the energy update when using entropy as an independent variable for a given evolutionary calculation step. Following Eq. 38, the term for the hydrogen-helium mixture is
Recent giant planet evolutionary calculations, such as those of Mankovich et al. (2016), and Mankovich & Fortney (2020), have approximated this term by differentiating the energy throughout the resolution of the profile (see Section 2.5 of Mankovich et al., 2016). Stellar evolution codes such as MESA (Paxton et al., 2011) have recently updated their fundamental energy equations to include explicitly the chemical potential (see Section 8 of Paxton et al., 2018). Along with the tables provided with this article, we also provide a means of obtaining this term for any , regardless of the interior profile.
7 One Jupiter Mass model examples
As an example of the use of these EOS tables, we employ the evolutionary code APPLE (Sur et al., 2024) to compute a pair of one MJ evolutionary models, with and without helium rain. We assume here the CD21 equation of state and the Lorenzen et al. (2011) miscibility curves, shifted by K131313Although this model is not designed to fit Jupiter in detail, such a shift is required to maintain a flat 3 solar metallicity envelope to closely match the observed abundance (Von Zahn et al., 1998).. These models do not purport to be fits to Jupiter, as a full and detailed study with APPLE is the subject of upcoming work (Sur et al. 2024b, in prep). The APPLE code includes, as best we currently can, helium rain. In it, we employ a scale parameter (, here set equal to cm) that governs the magnitude of an settling/advective term, while conserving mass and properly accounting for the gravitational energy release. The parameter results in an equilibrium helium “rain cloud” that is quasi-exponentially distributed in space (similar to models of condensation clouds, but consistent with the miscibility curves).
Figure 9 shows the thermal profile evolution as a function of pressure and mass on the left column, and the evolution of the entropy profiles and helium profiles on the right column. These models were computed using a flat water abundance profile at 3 times solar throughout the envelope, with a total metallicity mass of M⊕ for both models, assuming M⊕ of ice/rock/iron mixtures in a compact core. We used atmospheric boundary conditions that included solar irradiation and ammonia clouds as computed by Chen et al. (2023). While these models are merely test examples of the use of our tables in an evolutionary code, they do still incorporate much of the basic physics necessary to address gas giant evolution. More sophisticated models using APPLE will be forthcoming.
8 Conclusion
With this paper, and its associated repository, we provide a more centralized and organized way to access the crucial hydrogen-helium-metallicity mixture equations of state, accompanied by the various derivatives necessary to calculate heat transport and different convection criteria. The EOSes and derivatives are recalculated in the range and , a range that is comparable to each of the EOS tables that we have adapted. Importantly, we calculate the pseudo-chemical potential (Eq. 39) so that it is readily available for energy calculations for any helium mixture.
Our primary purpose is to provide curated access to equation of state and miscibility data and thermodynamic derivatives presently scattered in the literature. In addition, though the standard Schwarzchild and Ledoux conditions discussed in the stellar and planetary evolution literature (Kippenhahn & Weigert, 1990; Paxton et al., 2011; Gabriel et al., 2014; Mihalas, 1978) use temperature and composition as primary variables and refer centrally to and , the alternate set of variables (entropy and composition) have virtues (see Sur et al., 2024) and in this paper have provided the corresponding formulae.
The reader should note that the effects of radiative/conductive and composition diffusion on the Ledoux condition and the various approaches to doubly-diffusive instabilities (“semi-convection”) are discussed in many works (see, e.g., Kippenhahn & Weigert (1990), Paxton et al. (2011), Langer et al. (1983), and Tremblin et al. (2019)) and will not be repeated here. On doubly-diffusive instabilities, the group of Pascale Garaud has generated a series of important papers (Rosenblum et al., 2011; Traxler et al., 2011; Mirouh et al., 2012; Wood et al., 2013; Brown et al., 2013; Moll et al., 2016; Tulekeyev et al., 2024) in the giant planet context. While a full discussion of these additional issues is beyond the scope of this paper, mention should be made of the complicating effects of helium rain. The standard Ledoux condition doesn’t explicitly incorporate the latent heat of miscibility, nor the issue of the growth and advective settling of helium droplets in a gravitational field. Importantly, the latent heat of hydrogen/helium immiscibility for the various miscibility models is not provided in the papers that perform detailed miscibility calculations. As a result, it is premature to delve into this topic in detail until such data from modern calculations are available. We note that the latent heat effect is likely much smaller than the heating per helium nucleus provided by the settling rain itself in a gravitational field, and the latter effect is properly handled in extant giant planet evolution codes (e.g., Sur et al., 2024).
In addition, a study of the instantaneous effect of helium droplet settling on a generalized Ledoux condition must acknowledge that the regions of the giant planet where miscibility emerges are likely already convective and that they cool into the miscibility curve. This implies that there is already upwelling and mixing in the region for which there is incipient phase separation. This hydrodynamics is not well addressed by the standard linear stability gedanken experiment that leads to the Ledoux condition, or its standard generalizations. Moreover, mixing-length theory, the standard context of spherical evolutionary codes, is not well suited to this multi-phase, multi-dimensional physics. Clearly, this highlights the need for a more sophisticated approach going forward to the interesting physics of helium rain in planet evolutionary models.
In summary, this work is designed to facilitate the prompt calculation of the thermodynamic quantities and their derivatives required for updated giant planet and exoplanet modeling in the Juno and Cassini eras and to help new groups of modelers join this exciting field.
Data Availability
All data are available on Zenodo under an open-source Creative Commons Attribution license: https://doi.org/10.5281/zenodo.10659249 (catalog doi:10.5281/zenodo.10659249).
Appendix A Composition gradient term with entropy gradients
Recall that with a composition gradient the criterion for stability against convection is Eq. 22. Using the definition in Eq. 23,
Recognizing that
in Eq. 23 can be written in terms of two derivatives of the entropy with respect to the helium mass fraction Y,
where the two terms can be assigned the names and ,
such that and as discussed in §6. Using the pressure scale height, these terms can be written in terms of entropy and composition profile gradients,
Appendix B The Brunt Väisälä frequency in terms of entropy and composition
Let so that
implying that
Using the relation on the right-hand side of Eq. 34,
which becomes
Using
and converting to non-logarithmic derivatives, this becomes
which is the expression found in Lattimer & Mazurek (1981). Using the relation
the Brunt Väisälä frequency (squared) can now be written as